Mark's soccer team played 12 games. The ratio of games won to games lost was 2:1. To the nearest whole percent, what percent of the games did his team lose?
Answer by stanbon(48568)

2:1 is the same as 2x:x where x is number of games lost.
--------
2x + x =12
3x = 12
x = 4 (number of games lost)
2x= 8 (number of games won)
------
Percent lost = 4/12 = 1/3 = 33.3333%
====================
Cheers,
Stan H.